# Dynamic Hybrid Resource Utilisation and MCS-based Intelligent Layering

### v1

@@ Abstract
The coexistence of heterogeneous service classes in 5G—Enhanced Mobile Broadband (eMBB), Ultra-Reliable Low Latency Communication (URLLC), and Massive Machine-Type Communication (mMTC)—poses major challenges for meeting diverse Quality-of-Service (QoS) requirements under limited spectrum and power resources. Existing radio access network (RAN) slicing schemes typically optimise isolated layers or objectives, lacking physical-layer realism, slot-level adaptability, and interpretable per-slice performance metrics. This paper presents a joint optimisation framework that integrates Dynamic Hybrid Resource Utilisation with MCS-Based Intelligent Layering, formulated as a mixed-integer linear program (MILP) that jointly allocates bandwidth, power, and modulation and coding scheme (MCS) indices per slice. The model incorporates finite blocklength effects, channel misreporting, and correlated fading to ensure realistic operation. Two modes are implemented: a Baseline Mode ensuring resource-efficient QoS feasibility, and an Ideal-Chaser Mode minimising deviation from ideal per-slice rates. Simulation results show that the proposed approach achieves energy efficiencies above $10^{7}$ kb/J in Baseline Mode and sub-millisecond latency with near-ideal throughput in Ideal-Chaser Mode, outperforming recent optimisation and learning-based methods in delay, fairness, and reliability. The framework provides a unified, interpretable, and computationally tractable solution for dynamic cross-layer resource management in 5G and beyond networks.
@@ I Introduction
To address these challenges, this paper proposes Dynamic Hybrid Resource Utilisation and MCS-based Intelligent Layering, a slot-level MILP framework for the joint allocation of bandwidth, power, and MCS indices under realistic physical-layer conditions, including finite blocklength shaping, channel misreporting, and temporally correlated fading. The framework operates in two regimes: (i) a baseline mode that ensures strict QoS satisfaction with minimal resource usage, and (ii) an ideal chaser mode that tracks target per-slice rates by leveraging full bandwidth and power budgets.
Unlike deep reinforcement learning (DRL) Wbased approaches, which rely on stochastic policy exploration and may violate hard QoS or latency constraints during training, the proposed MILP formulation offers deterministic optimality, interpretability, and feasibility guarantees at each optimisation instance. This ensures predictable and transparent control behaviour, which is essential for latency-critical or reliability-sensitive network slices.
@@ II Literature Review
Chen et al. [chen2023radiosaber] proposed RadioSaber, a two-tier channel-aware scheduler combining inter-slice and intra-slice resource allocation, achieving throughput gains of 17–72% while maintaining fairness. Nevertheless, the model omits modulation and coding scheme (MCS) adaptation and does not incorporate finite blocklength effects. Pramanik et al. [pramanik2024cers] introduced CERS, a multi-objective optimizer that balances latency and cost, but it lacks fading dynamics and real-time MCS adaptation. Ejaz and Choudhury [ejaz2025milpsurvey] surveyed linear, integer, and mixed-integer programming solutions for RAN resource allocation, emphasising tractability issues and limited physical-layer realism.
Korrai et al. [korrai2023energyefficient] investigated mmWave RAN slicing via fractional programming for beam, resource, and power optimization, but the framework does not include KPI-level tracking or slice-specific MCS adaptation. Grings et al. [grings2025nasp] presented NASP, a hierarchical orchestration platform that reduces URLLC session setup latency by 93%, yet lacks link-layer scheduling features such as per-slice MCS adaptation.
@@
At the application layer, Balasingam et al. [balasingam2024zipper] introduced Zipper, a predictive control paradigm offering per-application throughput and latency assurances with significantly lower tail latency. However, the model lacks PHY-layer granularity and per-slot adaptation. Eskandari et al. [eskandari2025drl] developed an O-RAN aligned DRL framework addressing URLLC/eMBB trade-offs under service-level agreement (SLA) constraints but did not include per-slice bandwidth or MCS selection. Lorincz et al. [lorincz2024energy] surveyed energy-efficiency KPIs across 5G slices, providing valuable benchmarking metrics, though without coupling to dynamic physical-layer resource management.
@@
In summary, the proposed Dynamic Hybrid Resource Utilisation and MCS-based Intelligent Layering framework differs from existing optimisation and DRL-based approaches by offering a unified, physically accurate, and analytically interpretable control structure. It ensures deterministic QoS compliance, supports both resource-frugal and ideal-tracking operation, and provides a transparent mechanism to balance efficiency, fairness, and performance—an aspect largely unexplored in current 5G RAN slicing literature.

### v2

@@ Abstract
The coexistence of heterogeneous service classes in 5G Enhanced Mobile Broadband (eMBB), Ultra-Reliable Low Latency Communication (URLLC), and Massive Machine-Type Communication (mMTC) poses major challenges for meeting diverse Quality-of-Service (QoS) requirements under limited spectrum and power resources. Existing radio access network (RAN) slicing schemes typically optimise isolated layers or objectives, lacking physical-layer realism, slot-level adaptability, and interpretable per-slice performance metrics. This paper presents a joint optimisation framework that integrates Dynamic Hybrid Resource Utilisation with MCS-Based Intelligent Layering, formulated as a mixed-integer linear program (MILP) that jointly allocates bandwidth, power, and modulation and coding scheme (MCS) indices per slice. The model incorporates finite blocklength effects, channel misreporting, and correlated fading to ensure realistic operation. Two modes are implemented: a Baseline Mode that ensures resource-efficient QoS feasibility, and an Ideal-Chaser Mode that minimises deviation from ideal per-slice rates. Simulation results show that the proposed approach achieves energy efficiencies above $10^{7}$ kb/J in Baseline Mode and sub-millisecond latency with near-ideal throughput in Ideal-Chaser Mode, outperforming recent optimisation and learning-based methods in delay, fairness, and reliability. The framework provides a unified, interpretable, and computationally tractable solution for dynamic cross-layer resource management in 5G and beyond networks.
@@ I Introduction
To address these challenges, this paper proposes a Dynamic Hybrid Resource Utilisation and MCS-based Intelligent Layering (DHRI) framework, a slot-level MILP model for the joint allocation of bandwidth, power, and MCS indices under realistic physical-layer conditions, including finite blocklength shaping, channel misreporting, and temporally correlated fading. The framework operates in two regimes: (i) a baseline mode that ensures strict QoS satisfaction with minimal resource usage, and (ii) an ideal chaser mode that tracks target per-slice rates by leveraging full bandwidth and power budgets.
Unlike deep reinforcement learning (DRL) W-based approaches, which rely on stochastic policy exploration and may violate hard QoS or latency constraints during training, the proposed MILP formulation offers deterministic optimality, interpretability, and feasibility guarantees at each optimisation instance. This ensures predictable and transparent control behaviour, which is essential for latency-critical or reliability-sensitive network slices.
@@ II Literature Review
Chen et al. [chen2023radiosaber] proposed RadioSaber, a two-tier channel-aware scheduler combining inter-slice and intra-slice resource allocation, achieving throughput gains of 17–72% while maintaining fairness. Nevertheless, the model omits modulation and coding scheme (MCS) adaptation and does not incorporate finite blocklength effects. Pramanik et al. [pramanik2024cers] introduced CERS, a multi-objective optimiser that balances latency and cost, but it lacks fading dynamics and real-time MCS adaptation. Ejaz and Choudhury [ejaz2025milpsurvey] surveyed linear, integer, and mixed-integer programming solutions for RAN resource allocation, emphasising tractability issues and limited physical-layer realism.
Korrai et al. [korrai2023energyefficient] investigated mmWave RAN slicing via fractional programming for beam, resource, and power optimisation, but the framework does not include KPI-level tracking or slice-specific MCS adaptation. Grings et al. [grings2025nasp] presented NASP, a hierarchical orchestration platform that reduces URLLC session setup latency by 93%, yet lacks link-layer scheduling features such as per-slice MCS adaptation.
@@
At the application layer, Balasingam et al. [balasingam2024zipper] introduced Zipper, a predictive control paradigm offering per-application throughput and latency assurances with significantly lower tail latency. However, the model lacks PHY-layer granularity and per-slot adaptation. Eskandari et al. [eskandari2025drl] developed an O-RAN aligned DRL framework addressing URLLC/eMBB trade-offs under service-level agreement (SLA) constraints, but did not include per-slice bandwidth or MCS selection. Lorincz et al. [lorincz2024energy] surveyed energy-efficiency KPIs across 5G slices, providing valuable benchmarking metrics, though without coupling to dynamic physical-layer resource management.
@@
In summary, the proposed Dynamic Hybrid Resource Utilisation and MCS-based Intelligent Layering framework differs from existing optimisation and DRL-based approaches by offering a unified, physically accurate, and analytically interpretable control structure. It ensures deterministic QoS compliance, supports both resource-frugal and ideal-tracking operation, and provides a transparent mechanism to balance efficiency, fairness, and performance, an aspect largely unexplored in current 5G RAN slicing literature.